OpenStudy (anonymous):

write an operation in slope intercept form through (-4,1) with a slope 1/2

OpenStudy (anonymous):

Well.. Not sure I have a nicer way to explain it. We know that the slope is 1/2, so by the slope intercept form: \[ y = mx + b \] we can say \[ m = \frac{1}{2} \] Now to find b, we say this \[ y = \frac{1}{2}x + b \\ \text{For (-4,1): }\quad x = -4 \quad \text{and} \quad y = 1 \\ 1 = \frac{1}{2} \cdot (-4) + b \\ 1 = -2 + b \\ b = 3 \] So at last: \[ y = \frac{1}{2}x + 3 \]